-VP use loplace transform to solve y-ayt y = f(t), 9 (6)=1, 96)=1 f(t) o tas It- 5t25 you may use partial de comp but must
Answer #1

3: ca r Fle Solutio -sinat Page No Date yll-2y + y = f(t) pages y(o)=1 / Y(o)=1 -(1) Fito $ tx5 t-5 t 7,5 L {F(t)} -st fe F(Page No Date Padez taking laplace transform on bs. OF C Is y l1-2ylty ] = LY FLL) for L5.4 (t)} = 415) we have L&9 113 = 5² gDate. Page 3 scs-12 At B S S-1 (S-1) 2 1 A CS-1)2 + Bg (S-1) tcs Put 5-0 A = 1 put s -1 Рcu s=2 1 A + 2B tas 1 1 + 2 B +2 2BPage No. Date Pades 1{ bas-132 1 6-1,23 +22 It { / -3 11/4 -tet +2+t -zet takina invease laplaie transform on(1) -s) et-s +